### v1

# Sneak Path Current Modeling in Memristor Crossbar Arrays for Analog In-Memory Computing

Abstract
Memristor crossbar arrays have emerged as a key component for next-generation non-volatile memories, artificial neural networks, and analog in-memory computing (IMC) systems. By minimizing data transfer between the processor and memory, they offer substantial energy savings. However, a major design challenge in memristor crossbar arrays is the presence of sneak path currents, which degrade electrical performance, reduce noise margins, and limit reliable operations. This work presents a closed-form analytical framework based on IMEC A14 (1.4 nm) Technology for accurately estimating sneak path currents in memristor crossbar arrays. The proposed model captures the interdependence of key design parameters in memristor crossbar arrays, including array size, ON/OFF ratio of memristors, read voltage, and interconnect conditions, through mathematically derived relationships. It supports various practical configurations, such as different data patterns and connection strategies, enabling rapid and comprehensive sneak path current modeling. The sensitivity analysis includes how design parameters influence sneak path current and noise margin loss, underscoring the trade-offs involved in scaling crossbar arrays. Validation through SPICE simulations shows that the model achieves an error of less than 10.9% while being up to 4784 times faster than full circuit simulations. This analytical framework offers a powerful tool for quantitative assessment and pre-design/real-time optimization of memristor-based analog in-memory computing (IMC) architectures.
I Introduction
MEMRISTOR, widely recognized as the fourth fundamental circuit element, has become central to the development of next-generation non-volatile memory technologies and emerging analog in-memory computing (IMC) systems, addressing the scaling and efficiency challenges of traditional CMOS-based memories [chua2003memristor, pourbakhsh2016sizing, gong2012clock, edstrom2019data, gong2014variation, chen2018viewer, khan2023pawn]. The primary architectural manifestation of memristor-based memory is the high-density passive crossbar array, where each cell is located at the intersection of perpendicular wordlines and bitlines [vourkas2015memristive, shin2011analysis]. This structure provides exceptional storage density since each cell can be scaled to the nanometer regime without requiring a transistor [kim2012functional, jo2009high], thereby offering higher bit-per-area ratios than conventional memories [lastras2018resistive]. Combined with its non-volatility and fast read capability, the memristor crossbar helps bridge the performance gap in modern memory hierarchies [uppaluru2025carbon, 6730965, oli2022stuck, oli2022reliability]. Additionally, memristor plays a crucial role in analog computing by enabling fast and direct hardware implementation of vector matrix multiplication in-memory, such as in a neuromorphic computing system. Memristors can perform computation intrinsically through Ohm’s and Kirchhoff’s laws when placed in crossbar arrays. The output current produced is in the form of weighted summations of input voltage and stored conductance of the memristor. This characteristic of analog IMC results in highly parallel, energy-efficient acceleration for signal processing and Artificial Intelligence (AI) tasks, as it removes the need for frequent data transfer between memory and processors [hu2018memristor].
However, a key challenge in fully realizing the advantages of passive memristor crossbar arrays is the sneak path current [vontobel2009writing]. Sneak paths are unwanted parallel current routes that occur during read operations, diverting current away from the target cell. In unoptimized crossbars, sneak path currents can dominate the total current by exceeding 50%, leading to substantial read errors and increased power consumption [demin2020sneak]. The absence of gating elements and the fully connected structure of the crossbar cause currents to flow preferentially through low-resistance cells [cassuto2013sneak], inevitably creating a parallel resistance network that distorts readouts and causes high-resistance state (HRS) to be misinterpreted as low-resistance state (LRS) [gul2019addressing]. Moreover, the magnitude and distribution of these parasitic currents are highly data-dependent, varying dynamically with the stored information [tang2018comprehensive] or data patterns. Analyzing sneak path currents during the pre-design and ongoing design phases is crucial, as their data-dependent variability poses significant reliability challenges [liang2010size]. Without precise estimation of these parasitic currents, it becomes difficult to ensure stable performance in terms of noise margin, functional yield, and maximum achievable array size [chen2013comprehensive].
Some research has focused on mitigating this issue through various strategies. Architectural approaches often incorporate gating elements, such as diodes, transistors, or additional memristors, to suppress unwanted current flow, although at the cost of reduced array density. Alternatively, gateless techniques leverage multi-point read schemes, coding strategies, or logical operations to achieve sneak path free reading [zidan2013memristor, joshi2021sneak, manem2010design, linn2010complementary, shi2020research, kannan2013sneak]. However, the above studies only limit to the individual aspects such as data patterns, connection configurations, read voltages, or array sizes, rather than capture the combined influence of all key parameters. Consequently, a unified analytical framework that comprehensively models sneak path behavior across multiple interacting factors remains largely unexplored. In addition, recent Computer-Aided Design (CAD) frameworks such as CiMLoop [andrulis2024cimloop], DNN+NeuroSim [luo2019mlp+], MNSIM [zhu2023mnsim], and CrossSim [maram2023deep] have advanced the modeling and co-design of IMC architectures. For example, CiMLoop enables full-stack exploration from device to architecture; DNN+NeuroSim provides circuit to algorithm benchmarking for SRAM and emerging non-volatile memories; MNSIM supports hierarchical modeling of digital and analog IMC architectures; and CrossSim simulates ReRAM-based neural training under realistic device nonidealities. However, these tools neither consider sneak path effects nor employ purely passive memristor arrays; instead, they depend on selector or switching devices to mitigate interference, which substantially limits memory density. This highlights the need for a comprehensive analytical framework dedicated to modeling sneak path phenomena in high-density passive crossbar arrays.
A multi-parameter closed-form analytical model is proposed in this paper to provide a unified framework for evaluating the combined influence of critical design parameters, including device conductance (ON/OFF ratio of memristor), array size, read voltage, interconnect resistance, data patterns, and connection strategies. The proposed approach enables accurate pre-design/real-time estimations of key performance indicators such as functional yield, noise margin, and maximum feasible array size, thereby supporting the development of robust and predictable memristor-based architectures. Compared with conventional SPICE simulations, the proposed closed-form model offers several advantages: (1) Speed and efficiency: it provides immediate analytical results, whereas SPICE requires extensive computation, particularly for large arrays with numerous parasitic components; (2) Scalability: the closed-form model can efficiently handle analyses involving thousands or millions of devices, while SPICE simulations slow down significantly as system size increases; (3) Analytical clarity: unlike SPICE, which produces purely numerical outputs, the closed-form approach reveals explicit relationships among parameters such as resistance, voltage, and current; (4) Design optimization: it allows designers to quickly explore trade-offs in parameters like array size and bias voltage without iterative simulations; and (5) Integration with design tools: the derived equations can be easily incorporated into CAD workflows for rapid performance estimation during schematic, layout or system-level design, whereas SPICE models are less adaptable for high-level design automation. This paper will make the following contributions:
  • The dependence of the sneak path current in the memristor crossbar array is analyzed for important circuit parameters.
  • A closed-form model is proposed for the accurate and efficient estimation of the sneak path current in memristor arrays, and its accuracy and run-time performance are validated against SPICE simulations.
  • The sensitivity and noise margins of the sneak path current in the memristor crossbar array are also modeled and deeply analyzed.
  • The closed-form model is further compared with the state-of-the-art.
This paper is organized into the following sections. In Section II, models and parameters for the memristor, the interconnect network, and closed-form expressions are presented along with the connection mechanism in array configuration. The results and discussions for the sneak path current, noise margin, the effect of interconnect resistance, sensitivity analysis, and model validation are presented in Section III. Performance and property comparison with the state-of-the-art is provided in Section IV. Finally, this paper is concluded in Section V.
II Methodology
II-A Design Considerations
The foundation of this analysis is the nanoscale crossbar memristor model presented in [robinett2007demultiplexers] based on HP Technology, where the device current is defined in the following equation:
$I=K\sinh(\alpha V)$ (1)
where $\alpha$ indicates the degree of nonlinearity in the device. It indicates the sensitive change of device characteristics due to a unit change in voltage. I and V represent the current and voltage, respectively. The parameter $K$ is analogous to the electrical conductance of a linear resistance.
In this configuration, a TiO2 layer for a total thickness of the device, ${h}\leq 10\,\text{nm}$, is sandwiched between two platinum electrodes. The oxide layer has two constituent layers: a doped region containing oxygen vacancies and an undoped region composed of pure TiO2, as shown in Fig. 1 (a). This work is conducted using a parameter value of $\alpha$ = 3 [strukov2008missing] and current versus voltage characteristic for the memristor is shown in Fig. 1 (d) [robinett2007demultiplexers].
Figure 1: (a) Structure of the linear dopant drift memristor model. h represents the total thickness of the device. w represents the thickness of the doped region; (d) Current versus voltage characteristic for the memristor [robinett2007demultiplexers]. A $3\times 3$ memristive crossbar, where Vdd = V2 is the voltage applied in row 2, and current sensed from column 2: (b) the crossbar without sneak paths where the desired current path is marked in green line, (c) presence of a sneak path which is marked in red line, (e) FRC connection strategy with ‘All Ones’ (LRS) data pattern, (f) GRC connection strategy with ‘All Zeros’ (HRS) data pattern. Here, in GRC connection strategy, shorted row nodes are indicated as "nrs" and shorted column nodes are indicated as "ncs".
The current method presented in this paper employs the constant Kon and Koff as the ON and OFF states, respectively. To comprehensively assess the influence of sneak path currents, two representative data patterns are simulated: the ‘All Zeros’ pattern (best case scenario, minimum sneak path current effect, all the memristors are in the OFF state and the sneak paths are made of OFF resistances in series); and the ‘All Ones’ pattern (worst case scenario, maximum sneak path current effect). Grounded unselected rows and columns has been explored as a method to mitigate sneak path currents. Grounded rows and columns of unselected cells in an array provides alternating paths for leakage currents to flow instead of the sensing circuit. However, part of the sneak path current still flows through the sense circuit. Four terminal biasing configurations are analyzed: Floating Rows and Columns (FRC), Grounded Rows with Floating Columns (GRFC), Floating Rows with Grounded Columns (FRGC), and Grounded Rows and Columns (GRC), to examine the impact of different connection strategies on the overall array behavior. The grounded rows and columns are implemented using $0.001\,\Omega$ resistors (near zero), with the corresponding nodes labeled as node row shorted and node column shorted, respectively.
Fig. 1 presents an abstract depiction of the sneak path phenomenon in the crossbar array, along with schematic representations to illustrate the different combinations of data patterns and connection strategies. The device at the center of the array is the “target device," and the devices sharing a line with the target are called “half-selected devices". To read the “target device", the selected row is pulled up to Vdd and the selected column is grounded through a resistor. Sneak path occurs when current flows from Vdd to ground through unselected devices. Therefore, it affects the accuracy of the read operation by degrading the voltage and current delivered to the device. The crossbar interconnect network is modeled at the unit-cell level to accurately capture parasitic effects, incorporating line resistance, stray capacitance to ground, and fringing capacitance between adjacent lines. In this work, the interconnect design dimensions follow the IMEC A14 (1.4 nm) technology node, offering concrete guidance for both academic researchers and industry engineers [pei2025interconnect, pei2024ultra, liu2023cfet, pei2023technology, liu2024future, liu2023cfetted]. For the IMEC A14 interconnect, metal layers M1 and M2 are primarily employed for internal interconnect within the cell. Layers M1 and M3 exhibit identical geometries, as do M4 and M5. However, M3 can be used for local interconnect within subarrays, and M4 and M5 are for intermediate-level interconnect. Metal layers M6 and above maintain consistent geometries for intermediate- and global-level interconnects. Consequently, layers M3, M5, and M6 are selected to represent interconnects spanning from local to global scales and are investigated. Fig. 2 (a) presents the utility of metal layers (M1~M12) for different interconnects. The geometry of the Cu interconnect and the pitch (P) is shown in Fig. 2 (b). Here, W, T represent the width and thickness of Cu. The thickness is represented as the product of the width (W) and aspect ratio (ar), and the pitch (P) represents the distance between the centers of two adjacent interconnects. The unit-cell memristor array is shown in Fig. 2 (c) with the corresponding values of W, T, and P. The selected interconnect metal layers provide different parasitic resistance Rline and capacitance Cline. The corresponding ranges of interconnect parameters are listed in Table I.
TABLE I: Interconnect Geometry, Resistance, and Capacitance
Metal W(nm) P(nm) T(nm) $R_{\text{line}}$($\Omega$) $C_{\text{line}}$(fF) M3 16 28 49 3.122 3.871e-3 M5 16 28 28 5.869 2.871e-3 M6 40 80 80 0.7396 1.020e-2
TABLE II: Boundary Conditions for Closed-form Models
Para. Boundary Para. Boundary Para. Boundary Size 44×44:6464×6464 $K_{on}$ 1e-9:1e-7 $V_{dd}$ 1V:3V
Figure 2: (a) Metal layers M1~M12 are used for interconnect within the cell; (b) Sketch of Cu for interconnect material; (c) Top view of a unit-cell memristor crossbar array for M5/M4 with IMEC A14.
Figure 3: Simulation results for sneak path current in terms of array sizes, data patterns (‘All Zeros’ and ‘All Ones’), FRC connection strategy, and line resistance M3 (Rline = $3.122\,\Omega$) where the read voltages are incremented along the rows and ON/OFF ratio is incremented along the columns of the subplots. The rise in sneak path current with respect to the increment of the ON/OFF ratio along the columns as Kon = [1e-9, 3e-8, 5e-8, 8e-8, 1e-7] for a specific read voltage: (a~e) Vdd = 1.0 V; (f~j) Vdd = 1.5 V; (k~o) Vdd = 2.0 V; (p~t) Vdd = 2.5 V; (u~y) Vdd = 3 V. Noticeably, for ‘All Ones’ pattern the sneak path current increases rapidly for the increments of read voltage, ON/OFF ratio, array size compared to the minimal increment of All Zeros data pattern.
TABLE III: Coefficients of Closed-form Expression for M3
Coeff. All 1s FRC All 1s GRFC All 1s FRGC All 1s GRC All 0s FRC All 0s GRFC All 0s FRGC All 0s GRC
$C_{1}$ -2.765766e-04 -5.271610e-04 -3.118053e-05 -3.422062e-05 -2.768951e-04 -5.173225e-04 6.728880e-08 7.209101e-08
$C_{2}$ -3.552098e-05 -7.224109e-04 -9.791516e-04 -9.441148e-04 -2.337750e-05 -7.911723e-05 -1.496197e-04 -1.496397e-04
$C_{3}$ 4.599539e-03 2.568702e-03 -3.942867e-03 -3.851644e-03 4.602914e-03 4.479660e-03 -8.406540e-04 -8.401264e-04
$C_{4}$ 1.722779e-02 3.588961e-02 -1.070919e-02 -9.972648e-03 1.747226e-02 4.457329e-02 -1.435455e-03 -1.437187e-03
$C_{5}$ -4.296973e-04 -9.912720e-03 -3.415353e-02 -3.305335e-02 -1.110902e-03 -1.103802e-03 -2.031960e-03 -2.030510e-03
$C_{6}$ -1.275372e-03 -2.229373e-02 -1.746779e-01 -1.730254e-01 -1.192496e-03 -3.191236e-03 -6.087109e-03 -6.086547e-03
$C_{7}$ 9.867175e-01 6.749072e-01 -2.319980e-02 1.489677e-02 -3.901237e-02 -3.450177e-02 -6.308356e-02 -6.303070e-02
$C_{8}$ -1.056307e-01 -1.838965e-01 -3.936425e-01 -3.919864e-01 -1.029701e-01 -1.556039e-01 -2.579691e-02 -2.578021e-02
$C_{9}$ 1.529703e+00 1.871899e+00 1.029943e+00 1.054458e+00 1.521661e+00 2.102373e+00 2.995661e+00 2.995596e+00
$C_{10}$ -1.154712e+00 -3.592643e+00 -8.795638e+00 -8.468067e+00 -2.441334e+01 -2.416504e+01 -2.428478e+01 -2.428423e+01
TABLE IV: Coefficients of Closed-form Expression for M5
Coeff. All 1s FRC All 1s GRFC All 1s FRGC All 1s GRC All 0s FRC All 0s GRFC All 0s FRGC All 0s GRC
$C_{1}$ -2.764303e-04 -5.369728e-04 -3.317372e-05 -3.607482e-05 -2.795334e-04 -5.162575e-04 1.109504e-06 1.100590e-06
$C_{2}$ -5.936502e-05 -1.011910e-03 -1.426496e-03 -1.405009e-03 -4.490970e-05 -1.323556e-04 -2.524449e-04 -2.524109e-04
$C_{3}$ 4.487461e-03 1.771752e-03 -5.595220e-03 -5.549250e-03 4.561308e-03 4.208091e-03 -1.425890e-03 -1.425737e-03
$C_{4}$ 1.695349e-02 3.218355e-02 -1.690781e-02 -1.636382e-02 1.732497e-02 4.394577e-02 -2.493809e-03 -2.492790e-03
$C_{5}$ -7.048080e-04 -1.162975e-02 -3.488875e-02 -3.399748e-02 -2.417057e-04 -1.852363e-03 -3.402301e-03 -3.402806e-03
$C_{6}$ -2.179162e-03 -2.706075e-02 -1.823875e-01 -1.810900e-01 -1.532339e-03 -5.523961e-03 -1.053107e-02 -1.053004e-02
$C_{7}$ 9.783429e-01 6.228614e-01 -3.333252e-02 -2.208753e-03 -5.741777e-03 -5.769302e-02 -1.051583e-01 -1.051793e-01
$C_{8}$ -1.081555e-01 -1.901170e-01 -3.993790e-01 -3.982103e-01 -1.062961e-01 -1.624248e-01 -4.035391e-02 -4.034864e-02
$C_{9}$ 1.524105e+00 1.817728e+00 9.227569e-01 9.431813e-01 1.527636e+00 2.088758e+00 2.976207e+00 2.976206e+00
$C_{10}$ -1.225103e+00 -4.002326e+00 -8.767657e+00 -8.499305e+00 -2.411070e+01 -2.436257e+01 -2.464774e+01 -2.464794e+01
TABLE V: Coefficients of Closed-form Expression for M6
Coeff. All 1s FRC All 1s GRFC All 1s FRGC All 1s GRC All 0s FRC All 0s GRFC All 0s FRGC All 0s GRC
$C_{1}$ -2.765330e-04 -5.136305e-04 -1.469021e-05 -1.697019e-05 -2.744689e-04 -5.168549e-04 -1.772653e-07 -1.834111e-07
$C_{2}$ -9.275605e-06 -3.903261e-04 -3.335456e-04 -2.664594e-04 -1.905130e-05 -2.198834e-05 -3.958725e-05 -3.958864e-05
$C_{3}$ 4.723620e-03 3.578035e-03 -1.398252e-03 -1.188272e-03 4.644600e-03 4.760412e-03 -2.236606e-04 -2.236878e-04
$C_{4}$ 1.751128e-02 3.976450e-02 -3.205943e-03 -2.113479e-03 1.729086e-02 4.514727e-02 -3.611497e-04 -3.607864e-04
$C_{5}$ -1.186960e-04 -7.747277e-03 -3.295507e-02 -3.143524e-02 -6.912710e-04 -2.288891e-04 -5.495305e-04 -5.487445e-04
$C_{6}$ -3.378056e-04 -1.659931e-02 -1.633178e-01 -1.609204e-01 -5.466582e-05 -7.376962e-04 -1.586451e-03 -1.588074e-03
$C_{7}$ 9.963053e-01 7.411125e-01 -5.549090e-03 4.635883e-02 -2.498576e-02 -6.894301e-03 -1.715344e-02 -1.712207e-02
$C_{8}$ -1.029360e-01 -1.751892e-01 -3.814526e-01 -3.785520e-01 -1.014793e-01 -1.482829e-01 -1.070043e-02 -1.069628e-02
$C_{9}$ 1.534969e+00 1.929448e+00 1.164724e+00 1.196345e+00 1.538061e+00 2.116129e+00 3.012977e+00 3.012930e+00
$C_{10}$ -1.073100e+00 -3.058583e+00 -8.790361e+00 -8.344802e+00 -2.429351e+01 -2.392658e+01 -2.388512e+01 -2.388481e+01
II-B Closed-form Model
Closed-form expressions for the sneak current are presented corresponding to different input variables, e.g., Size (Array Size), Kon (ON/OFF ratio), and Vdd (Read Voltage), as Equation (II-B). The boundaries for closed-form equation fitting are selected to encompass the full range of practical circuit biasing conditions and fabrication considerations. The corresponding ranges are listed in Table II. As shown in Equation (II-B), the corresponding coefficients are summarized in Tables IIIIVV.
$I_{sneak}=exp(C_{1}\cdot Size^{2}+C_{2}\cdot Size\cdot ln(K_{on})$
$+C_{3}\cdot Size\cdot V_{dd}+C_{4}\cdot Size+C_{5}\cdot ln(K_{on})^{2}$
$+C_{6}\cdot ln(K_{on})\cdot V_{dd}+C_{7}\cdot ln(K_{on})+C_{8}\cdot V_{dd}^{2}\textbf{}$
$+C_{9}\cdot V_{dd}+C_{10})$ (2)
TABLE VI: Sneak Path Current Model Validation: Simulation Results, Errors, and Normalized Run-time Improvement
Pattern Strategy Metal Size:8×8, Kon:3e-8, Vdd:1.5V Size:16×16, Kon:5e-8, Vdd:2V Size:32×32, Kon:8e-8, Vdd:2.5V
Simulation(A) Error Run-time Simulation(A) Error Run-time Simulation(A) Error Run-time
All 1s FRC M3 1.089e-07 -7.017% 1916× 3.896e-07 -6.257% 835× 1.312e-06 2.163% 2030×
All 1s FRC M5 1.089e-07 -6.895% 768× 3.895e-07 -6.213% 159× 1.307e-06 2.052% 4784×
All 1s FRC M6 1.089e-07 -7.151% 1132× 3.897e-07 -6.319% 1466× 1.316e-06 2.325% 1821×
All 1s GRFC M3 3.608e-07 -6.321% 1198× 1.866e-06 -7.207% 62× 8.210e-06 7.760% 1777×
All 1s GRFC M5 3.608e-07 -5.789% 1103× 1.861e-06 -7.205% 2048× 7.935e-06 7.308% 1464×
All 1s GRFC M6 3.609e-07 -6.993% 690× 1.870e-06 -7.405% 1355× 8.479e-06 8.717% 1023×
All 1s FRGC M3 1.312e-06 10.573% 730× 8.314e-06 -4.316% 542× 2.969e-05 -3.853% 1618×
All 1s FRGC M5 1.312e-06 10.857% 1378× 8.202e-06 -4.910% 137× 2.669e-05 -2.296% 1561×
All 1s FRGC M6 1.313e-06 10.043% 43× 8.415e-06 -3.978% 888× 3.336e-05 -4.605% 1876×
All 1s GRC M3 1.313e-06 10.260% 10× 8.331e-06 -4.263% 1024× 3.014e-05 -4.124% 1230×
All 1s GRC M5 1.312e-06 10.618% 172× 8.218e-06 -4.857% 779× 2.703e-05 -2.529% 1694×
All 1s GRC M6 1.313e-06 9.584% 1091× 8.433e-06 -3.967% 2426× 3.401e-05 -4.837% 1698×
All 0s FRC M3 3.630e-10 -6.942% 826× 7.800e-10 -6.381% 305× 1.640e-09 2.166% 1539×
All 0s FRC M5 3.630e-10 -7.075% 746× 7.800e-10 -6.337% 1392× 1.630e-09 2.753% 1646×
All 0s FRC M6 3.630e-10 -6.964% 214× 7.800e-10 -6.286% 1382× 1.650e-09 1.980% 1229×
All 0s GRFC M3 1.208e-09 -8.582% 123× 3.810e-09 -7.350% 1269× 1.162e-08 8.837% 1385×
All 0s GRFC M5 1.208e-09 -8.254% 977× 3.810e-09 -7.298% 1489× 1.151e-08 8.491% 1672×
All 0s GRFC M6 1.208e-09 -8.930% 684× 3.820e-09 -7.701% 319× 1.171e-08 9.336% 1873×
All 0s FRGC M3 4.499e-09 1.050% 965× 2.013e-08 0.573% 600× 8.839e-08 -1.321% 1827×
All 0s FRGC M5 4.499e-09 1.765% 1263× 2.011e-08 0.837% 1352× 8.686e-08 -2.087% 1814×
All 0s FRGC M6 4.500e-09 0.283% 774× 2.016e-08 0.167% 1615× 8.978e-08 -0.420% 1167×
All 0s GRC M3 4.499e-09 1.049% 806× 2.013e-08 0.571% 1176× 8.839e-08 -1.321% 1748×
All 0s GRC M5 4.499e-09 1.765% 954× 2.011e-08 0.837% 930× 8.686e-08 -2.085% 945×
All 0s GRC M6 4.500e-09 0.281% 1078× 2.016e-08 0.166% 1455× 8.978e-08 -0.421% 800×
III Results and Discussion
III-A Sneak Path Current Analysis
The sneak path effect is examined for different memristor crossbar array sizes; the analysis also includes a range of device parameters and data patterns. The magnitude of the sneak path current and the corresponding array size have a strong correlation. As the size of the array increases, the number of potential parallel sneak paths increases, leading to an increase in the current value of the sneak path for a given memristor resistance at a given input voltage value. Fig. 3 demonstrates the scalability challenge, where dense arrays face exponentially worse sneak path issues. The effect of the data pattern in the array is also shown in Fig. 3, where all memristors are in a LRS, creating the maximum number of low-resistance parallel paths. Consequently, the “All Ones" pattern consistently results in the highest sneak path current across all array sizes and parameter sets. Similarly, all memristors with a HRS present the highest overall resistance to current flow. As a result, the “All Zeros" pattern consistently yields the lowest sneak path currents in all tested conditions. Additionally, Fig. 3 illustrates how different device parameters can either mitigate or aggravate the sneak path problem. Considering the FRC connection strategy with M3 as the interconnect metal, the dependence of Kon value and Vdd value highly influences the sneak path current as shown in Fig. 3 (a-y). As the read voltage (Vdd) increases from 1 V to 3 V with a step size of 0.5 V, a significant increase in sneak path current can be observed in Fig. 3 (a, f, k, p, u). The Kon values are swept while keeping the Vdd fixed, and the trend of sneak path current is observed across five rows in Fig. 3. The first row indicates the sweeping of Kon with a linear increment for a fixed Vdd, while the next four rows have similar Kon sweeping but have different Vdd values across consecutive rows. As the ON/OFF ratio of memristor gets increased by having Koff set at 1e-10, and changing Kon to 1e-9, 3e-8, 5e-8, 8e-8, and 1e-7, the sneak path current also increases, which is observed in Fig. 3. Noticeably, the sneak path current is negligible for small array sizes, but increases significantly for large array sizes for both ‘All Ones’ and ‘All Zeros’ data patterns. However, for the ‘All Zeros’ pattern, the increase of sneak path current is significantly lower at a higher ON/OFF ratio as the resistances tend to become ineffective at a higher ON/OFF ratio.
III-B Noise Margin Analysis
The allowable voltage levels in a CMOS circuit for logic ZERO should be ideally VOL= GND, and for logic ONE, it should be VOH = Vdd. However, some variations in these values are acceptable under specific bounds, which are referred to as ‘Noise Margin’. The noise margin of a memristor cell is defined as the following equation:
$Margin_{noise}=V_{one}-V_{zero}$ (3)
where $V_{one}$ and $V_{zero}$ denote the voltages of the target cell in the low-resistance and high-resistance states, respectively. For a memristor crossbar array, the voltage values representing logic ONE and ZERO are highly affected due to the sneak path current. The noise margin of a target memristor cell is a strong indicator of whether the cell is operating effectively in an array. A large noise margin means the sensing circuit can easily identify the stored bits despite the sneak path effect and other disturbances. In contrast, a small noise margin usually means that the sensed voltages are too close to distinguish, which makes the cell unreliable and effectively nonfunctional in the array. The sneak path effect on noise margin is indicated by the difference between the sensed voltages corresponding to stored ZEROs and ONEs at the target cell.
For the analysis of the sneak path effect on the noise margin of the memristor array, a normalized parameter is defined [zidan2013memristor]:
$Margin_{noise}=\frac{Margin_{noise\_array}}{Margin_{noise\_device}}$ (4)
Here, Marginnoise_device represents the noise margin of a single device without any interference of the sneak path, as the single device does not counter the sneak path issue, and Marginnoise_array represents the voltage difference of the ON and OFF states of a target cell while being subjected to operation in a memristor array. The noise margin of a target cell in an array highly depends on the data pattern stored within memory. For our analysis, only the best case (All Zeros) and the worst case (All Ones) data patterns are introduced.
Taking M3 as an example, the simulation results in Fig. 4 demonstrate that at small array sizes the noise margins of both ‘All Ones’ and ‘All Zeros’ are almost similar at lower read voltage, e.g., Vdd= 1 V and smaller ON/OFF ratio, e.g., Kon= 1e-9. However, as the read voltage and Kon increase, the noise margin of larger arrays tends to decrease significantly. This is due to the increase of the sneak path current in larger arrays for high read voltage and high ON/OFF ratio. Finally, from the results, the added resistance of the sneak paths significantly narrows the noise margin and reduces the maximum possible size of a memristor array. It is the limiting factor for array scalability.
Figure 4: Simulation results for noise margin versus array sizes, data patterns (‘All Zeros’ and ‘All Ones’), FRC connection strategy, and line resistance M3 (Rline = $3.122\,\Omega$) where the read voltages are incremented along the rows and ON/OFF ratio is incremented along the columns of the subplots. The fall of noise margin for ‘All Ones’ data pattern with respect to the increment of the ON/OFF ratio along the columns as Kon = [1e-9, 3e-8, 5e-8, 8e-8, 1e-7] for a specific read voltage: (a~e) Vdd = 1.0 V; (f~j) Vdd = 1.5 V; (k~o) Vdd = 2.0 V; (p~t) Vdd = 2.5 V; (u~y) Vdd = 3 V. Noticeably, for ‘All Ones’ pattern, noise margin rapidly decreases with respect to the increments of read voltage, ON/OFF ratio, array size compared to approximately zero decrease for ‘All Zeros’ data pattern.
III-C Interconnect Resistance Effect on Sneak Path Current
[…]
Figure 5: (a) Sneak path current and (b) noise margin of the target cell versus array size for the interconnect layers using the IMEC A14 technology node. The array adopts ‘All Ones’ data pattern with FRC connection strategy. Kon is 1e-7. Vdd is 3 V.
Figure 6: Relative percentage change in sneak path current and noise margin as Kon and Vdd are varied. The array adopts ‘All Ones’ data pattern with FRC connection strategy using the IMEC A14 technology node with Cu M3 interconnect.
III-D Model Validation Against SPICE Simulation
To assess the accuracy and applicability of the closed-form model of the sneak path current, three sets of samples are selected for each parameter, and the resulting errors are analyzed. For validation, intermediate points are selected instead of boundary points to avoid boundary effects, as fitting constraints at the boundaries may produce artificially small or near-zero errors. Table VI presents the SPICE simulation results and model errors with high efficiency. All errors are below 10.9%, indicating that the proposed model accurately captures the sneak path current. In addition, under the same hardware environment, the proposed model achieves a run-time 10 to 4784 times faster than that of SPICE simulations. The hardware specifications are as follows: CPU Intel Core Ultra 7 256V, base frequency 2.2 GHz, 1 socket, 8 cores, 8 logical processors, and 16 GB of memory.
III-E Sensitivity Analysis
A detailed and systematic analysis of sneak path currents within passive memristor crossbar arrays is presented, thereby expanding the current research landscape regarding parameter coverage and the depth of analytical modeling. The parameters selected within this study aim to effectively connect circuit-level characteristics with fabrication-level considerations, thus providing a practical framework that integrates theoretical models with physical implementation. Sneak path currents are vital in evaluating the feasibility of array designs. Excessive sneak currents can significantly constrain the maximum achievable size of the array, decrease the noise margin of target cells, and hurt data integrity during read operations. Additionally, the ON/OFF resistance ratio of the devices plays an important role in influencing the severity of sneak paths: a lower contrast between the conductive and resistive states can mitigate undesired currents, whereas a higher ratio may intensify these currents. Moreover, this analysis explicitly considers the impact of interconnect line resistance, which becomes increasingly significant as devices and arrays continue to scale. In advanced technologies, line resistance contributes to voltage degradation along word lines and bit lines, thereby influencing write reliability and read stability. From the closed-form analysis, the observed errors result primarily from the limited number of sampled points in SPICE simulations. To balance run-time and model accuracy, the current number of sampling points is limited and array size is set from 4×4 to 64×64. Actually, the resulting errors could be further reduced by increasing the number of points or optimizing the model. Also, the close-form model has good scalability to be applicable for larger array size with the same methodology, which represents a promising direction for future work. Compared with full circuit-level simulations, the proposed model achieves comparable accuracy with substantially reduced time. The fast computing characteristics make the model suitable for large-scale array analysis and design optimization. To quantify the relative impact of array size on the noise margin and sneak path current, respectively, the size-to-margin sensitivity (Zn) and size-to-current sensitivity (Zi) for noise margin and sneak path current are defined in the following equations. A larger absolute value indicates a stronger sensitivity.
$Z_{n}=\frac{Margin_{noise\_64\times 64}-Margin_{noise\_4\times 4}}{Margin_{noise\_4\times 4}}$ (5)
$Z_{i}=\frac{I_{sneak\_64\times 64}-I_{sneak\_4\times 4}}{I_{sneak\_4\times 4}}$ (6)
Figure 7: Radar chart for sensitivity of sneak path current and noise margin to key design parameters.
Taking M3 as an example, Fig. 6 presents the simulation results to quantify the tradeoff among critical design parameters on Isneak and noise margin. The relative percentage change surface for both sneak current and noise margin is shown with respect to the change of Vdd and Kon values. In this work, the key input parameters across all connection strategies and data patterns are varied systematically, adjusting Vdd by a factor of (3V - 1V)/1V = 2, Kon for (1e-7-1e-9)/1e-9 = 99, and array size for (64×64 - 4×4)/(4×4) = 255. The resulting impact on sneak path current and noise margin is quantitatively ranked to assess parameter sensitivity. This analysis reveals how each parameter contributes to either aggravating or mitigating sneak path effects. Among the three, Vdd exhibits the highest sensitivity, while array size shows the lowest. Such pre-design evaluation enables early prediction of how expanding the conductance range or lowering the read voltage can support larger array sizes, and vice versa. The trend and comparison for sensitivity are shown in a radar chart, Fig. 7.
IV Comparison with State-of-The-Art
TABLE VII: Comparison with State-of-the-Art
Device Type [zidan2013memristor] [manem2010design] [joshi2021sneak] [datta2022unlocking] Current
Array size consideration $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Read voltage consideration $\surd$ × x $\surd$ $\surd$
ON/OFF ratio consideration $\surd$ × $\surd$ $\surd$ $\surd$
Data pattern consideration $\surd$ × $\surd$ × $\surd$
Connection strategy consideration $\surd$ $\surd$ × × $\surd$
Interconnect model consideration $\surd$ × × × $\surd$
Closed-form equation consideration × × $\surd$ × $\surd$
Accuracy consideration × × × × $\surd$
Table VII provides a thorough comparison between the proposed analytical framework and existing studies on memristor-based systems. While previous research has addressed specific design parameters that influence sneak path behavior, none have comprehensively examined the complete set of interrelated factors affecting array performance. Key parameters, including array size, connection strategy, ON/OFF ratio, interconnect resistance, and read voltage, are essential for accurately evaluating the reliability and scalability of passive memristive memory arrays. In contrast, this study presents a novel and effective analytical approach that integrates all these factors within a cohesive framework. The proposed closed-form models demonstrate minimal error across various validation conditions, thus facilitating precise pre-design and real-time estimations, as well as enhanced performance predictions for memristor-based crossbar systems.
V Conclusion
Sneak path currents significantly disrupt the reliable operation of passive memristor crossbar arrays. Therefore, estimating the sneak path effect during both the pre-design and ongoing design phases is essential for determining feasible array sizes and performance limitations. This work introduces an analytical framework for estimating sneak paths across a diverse range of design parameters and array configurations. The resulting closed-form models demonstrate the relationships of sneak path current with various parameters, revealing linear, exponential, and logarithmic dependencies under different biasing conditions. The proposed model is highly efficient, achieving an error of less than 10.9 % while reducing computation time by a factor of 10 to 4784. This enables accurate and efficient estimation of the impact of sneak paths. The accompanying sensitivity analysis further highlights how variations in design parameters affect both sneak path current and noise margin, providing essential guidance for balancing competing design trade-offs. Such detailed assessments offer designers valuable insights into scalability, functional yield, and performance constraints for future memristor-based analog IMC systems.
Acknowledgment
This work is supported in part by the National Science Foundation under grants 2510192, 2502054, 2428981, 2420994, 2427766, 2408064, 2247343, and 2218046. This material was also based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research (ASCR), under Award Number(s) DE-SC0025561.
[Uncaptioned image] Shah Zayed Riam received the B.Sc. and M.S. degrees in electrical engineering from Rajshahi University of Engineering and Technology (RUET), Bangladesh, and the University of Texas at Tyler, USA, respectively, in 2022 and 2024. He is currently pursuing a Ph.D. degree at the University of Alabama, Tuscaloosa, AL, USA. His research interests include memory design, flexible electronics, and electrochemical biosensors. His previous works received a runner-up award at WiSe/YP BIP 2023 and a nomination for the Best Paper Award at IEEE FLEPS 2023, Boston, MA, USA.
[Uncaptioned image] Zhenlin Pei (Graduate Student Member, IEEE) holds the M.S. and Ph.D. degrees in electrical engineering from Columbia University and the University of Texas at Arlington. During his Ph.D., he collaborated with the Interuniversity Microelectronics Centre to develop the Cacti++ framework, which bridges gaps in rapid EDA-based co-design and co-optimization from the transistor to the system level. He is a postdoctoral fellow in the department of electrical & computer engineering at the University of Alabama. He was a senior design engineer in the IP group for tapeout at Cadence for four years. Research interests include energy-efficient computing from hardware to software, with applications in AI through CAD/EDA & DTCO/STCO, emerging interconnects/beyond-CMOS/memory technologies, AI hardware with privacy, and neuromorphic systems.
[Uncaptioned image] Kyle Mooney (Graduate Student Member, IEEE) received his B.S. degree in computer engineering from the University of South Alabama in 2023 and is currently pursuing a Ph.D. in Electrical Engineering from the University of Alabama. His research interests include memory design, edge computing, and artificial intelligence.
[Uncaptioned image] Chenyun Pan (Senior Member, IEEE) received a B.S. in Microelectronics from Shanghai Jiao Tong University and a Ph.D. in ECE from Georgia Institute of Technology. He is an Associate Professor at the University of Texas at Arlington, and his research focuses on energy-efficient Boolean and non-Boolean computing systems using emerging technologies. He has published over 80 IEEE papers and received several awards, including the Research Spotlight Award from Georgia Institute of Technology, and an early career research award from the U.S. Department of Energy.
[Uncaptioned image] Na Gong (Senior Member, IEEE) received the Ph.D. degree in computer science and engineering from the State University of New York, Buffalo, in 2013. Currently, Dr. Gong is a professor in the Department of Electrical and Computer Engineering at the University of Alabama. Her research interests include power-efficient computing circuits and systems, memory optimization, AI hardware, and hardware privacy. She is the recipient of the best paper nomination from ISVLSI’19, best paper award from EIT’16, best paper nominations from ISQED’16 and ISLPED’16.
[Uncaptioned image] Jinhui Wang (Senior Member, IEEE) is currently a Full Professor and Larry Drummond Endowed Chair with the Department of Electrical and Computer Engineering at the University of Alabama, Tuscaloosa, AL, USA. His research interests include: (1) VLSI System, Digital and Mixed-Signal Integrated Circuit (IC) Design, 3D and 2.5D IC Design, and Emerging Memory; (2) AI Hardware Design, Post/Beyond CMOS Device, such as Memristors, Based Neuromorphic Computing System; and (3) Post/Beyond CMOS Devices Enabled Cybersecurity and Internet of Things (IoT) Systems. He has published over 200 refereed journal/conference papers and book chapters as well as 31 patents in the area of emerging semiconductor technologies. His previous work has received the Best Paper Award/Nomination at DATE 2021, ISVLSI 2019, ISLPED 2016, ISQED 2016, and EIT 2016.